# Invariants of Finite Orthogonal Groups of Plus Type in Odd Characteristic

### v2

@@ 4. Formulating the Main Theorem
In the following, for $i<2m$, we identify $R_{i}$ and $\Phi_{i,m}(R_{i})$. Recall that $\psi:S_{m}\to S_{m}[t]$ is the algebra homomorphism determined by $\psi(v)=v^{q}-vt^{q-1}$ for $v$ homogeneous of degree one. Note that $\psi$ followed by evaluating $t$ at $x_{1}$ gives $\psi_{1}:S_{m}\to S_{m}$. Also note that $\psi(\xi_{0})=\xi_{0}^{q}-\xi_{1}t^{q-1}+\xi_{0}t^{2(q-1)}$, $\psi(\xi_{1})=\xi_{1}^{q}-\xi_{2}t^{q-1}-2\xi_{0}^{q}t^{q(q-1)}+\xi_{1}t^{(q+1)(q-1)}$, and $\psi(\xi_{i})=\xi_{i}^{q}-\xi_{i+1}t^{q-1}-\xi_{i-1}^{q}t^{q^{i}(q-1)}+\xi_{i}t^{(q^{i}+1)(q-1)}$ for $i>1$. Therefore $\psi$ restricts to a map from $R_{n-2}$ to $R_{n-1}[t]$.
For $f\in R_{n-1}$, define ${R\mbox{-}\deg}(f)$ to be the degree of $f$ as a polynomial in the variables $\xi_{0},\ldots,\xi_{n-1}$. We will use the corresponding weighted grevlex order on $R_{n-1}$.
@@
Define $\nu:R_{n-1}\to\mathbb{Z}\cup\{\infty\}$ by $\nu(0)=\infty$ and, for $f\not=0$, $\nu(f)$ is the minimum value of ${R\mbox{-}\deg}$ on the non-zero terms of $f$. For example, $\nu(\xi_{1}^{q})=q$ and $\nu(\xi_{1}^{q}-\xi_{2}\xi_{1}^{(q-1)/2})=(q+1)/2$. Note that if $\nu(f)=\ell$, then $f\in(R_{n-1})_{+}^{\ell}\setminus(R_{n-1})_{+}^{\ell+1}$. It is easy to verify that for $f,h\in R_{n-1}$, we have $\nu(fh)=\nu(f)\nu(h)$ and $\nu(f+h)\geq{\rm min}\{\nu(f),\nu(h)\}$. It follows from [25, Ch IV, §9, Theorem 14], that defining $\nu(f/h)=\nu(f)-\nu(h)$ extends $\nu$ to a discrete valuation on $\mathcal{F}(R_{n-1})$. Note that the associated Discrete Valuation Ring, $\{a\in\mathcal{F}(R_{n-1})\mid\nu(a)\geq 0\}$, properly contains $R_{n-1}$, i.e., $R_{n-1}$ is not the associated Discrete Valuation Ring.
@@
For $f\in R_{n-1}$, we have $\nu(\mathcal{P}^{i}(f))\geq\nu(f)$.
For $f\in R_{n-1}$, define ${S\mbox{-}\deg}(f)$ to be the degree of $f$ as an element of $S_{m}$. For example, ${S\mbox{-}\deg}(\xi_{i})=q^{i}+1$ while ${R\mbox{-}\deg}(\xi_{i})=1$.
@@
Suppose $\beta$ is a monomial in $R_{n-1}$ with ${S\mbox{-}\deg}(\beta)={S\mbox{-}\deg}(\xi_{i}^{b}\xi_{i+1}^{a})$. If $\nu(\beta)>\nu(\xi_{i}^{b}\xi_{i+1}^{a})$ then $\operatorname{LT}(\beta)<\operatorname{LT}(\xi_{i}^{b}\xi_{i+1}^{a})$ using the weighted grevlex order.
@@
It is sufficient to show that $\beta$ is in the ideal $\langle\xi_{0},\ldots,\xi_{i-1},\xi_{i}^{b+1}\rangle\subset R_{n-1}$. Suppose, by way of contradiction, that this is not the case. Since $\nu(\beta)>b+a$, $\beta$ has at least $a+b+1$ factors. Since $\beta$ is not in the ideal $\langle\xi_{0},\ldots,\xi_{i-1},\xi_{i}^{b+1}\rangle$, at least $a+1$ of these factors must be greater than $\xi_{i}$. The smallest weighted degree of such a monomial is
$b(q^{i}+1)+(a+1)(q^{i+1}+1)>b(q^{i}+1)+a(q^{i+1}+1)={S\mbox{-}\deg}(\xi_{i}^{b}\xi_{i+1}^{a}),$
@@
$M_{m}:=\begin{pmatrix}\xi_{n-1}&\xi_{n-2}&&\cdots&\xi_{m-1}\\ \xi_{n-2}^{q}&\xi_{n-3}^{q}&&\cdots&\xi_{m-2}^{q}\\ &&\vdots&&\\ \xi_{m}^{q^{m-1}}&\xi_{m-1}^{q^{m-1}}&&\cdots&\xi_{0}^{q^{m-1}}\end{pmatrix}$
and let $M(i,m)$ denote the minor formed by removing column $i+1$ from $M_{m}$. Note that $\nu(M(i,m))=1+q+\cdots+q^{m-1}$. It follows from Lemma 6.1 that $\nu(u_{2})=q+1$ and $u_{2}=M(0,2)+\delta_{0,2}$ with $\nu(\delta_{0,2})=2q$. It follows from Lemma 6.4 that $u_{2}d_{1,2}=M(1,2)+\delta_{1,2}$ with $\nu(\delta_{1,2})=2q$ and $u_{2}d_{2,2}=M(2,2)+\delta_{2,2}$ with $\nu(\delta_{2,2})=q^{2}+1$.
@@
Note that $\nu(m_{j,k})=q^{j-1}$. For $k>1$, it follows from Lemma 3.5 that $\mathcal{P}^{q^{n-k}}(m_{j,k})=m_{j,k-1}$, and $\nu(\mathcal{P}^{\ell}(m_{j,k}))>q^{j-1}$ unless $\ell=0$ or $\ell=q^{n-k}$. Therefore, since $e(i,m)=q^{n-2}+q^{n-3}+\cdots+q^{n-i-1}$, we have
@@
We will show below that $u_{m}d_{i,m}=M(i,m)+\delta_{i,m}$ with $\nu(\delta_{i,m})>\nu(M(i,m))=1+q+\cdots+q^{m-1}$.
Finally we are ready to formulate the main theorem. Recall from Theorem 3.11 the homogeneous system of parameters
=$\mathcal{H}=\{\xi_{0},\ldots,\xi_{m-1},d_{1,m},\ldots,d_{m,m}\}$
and let $\mathcal{B}$ denote the set of monomial factors of $\prod_{i=m}^{n-2}\xi_{i}^{q^{n-i}-1}$.
@@
Suppose $m>1$.
(a) $u_{m}\in R_{n-2}$, $\ker(\Phi_{n-2,m-1})=u_{m}R_{n-2}$, $u_{m}d_{i,m}\in R_{n-1}$ and $d_{i,m}\in S_{m}^{G_{m}}$. Furthermore, interpreting $u_{m-1}$ and $u_{m-1}d_{i,m-1}$ as elements of $R_{n-3}\subset S_{m}^{G_{m}}$, we have
$u_{m}=(\xi_{n-2}+c_{n-2,m})u_{m-1}^{q}+\sum_{i=1}^{m-1}(-1)^{i}(\xi_{n-2-i}+c_{n-2-i,m})(u_{m-1}d_{i,m-1})^{q}$
with $c_{j,m}\in R_{j-1}$, $u_{m}d_{i,m}-(u_{m-1}d_{i-1,m-1})^{q}\xi_{n-1}\in R_{n-2}$, $\mathcal{P}^{i}(u_{m})=0$ for $0<i<q^{m-1}$ and $\mathcal{P}^{i}(u_{m})/u_{m}\in R_{n-2}$ for $q^{m-1}\leq i<q^{n-2}$. For $i=\ell q^{n-2}+r$ with $0\leq\ell<q$ and $0\leq r<q^{n-2}$, $u_{m}^{\ell+1}\mathcal{P}^{i}(d_{j,m})\in R_{n-1}$ with $\xi_{n-1}$-degree at most $\ell+1$. If $\ell=q-1$ and $i\not\equiv_{(q)}0$ then $u_{m}^{q}\mathcal{P}^{i}(d_{j,m})$ has $\xi_{n-1}$-degree at most $q-1$.
(b) Using the weighted grevlex order on $R_{n-1}$,
$\operatorname{LT}(u_{m})=(-1)^{\lfloor m/2\rfloor}\xi_{m-1}^{q^{m-1}+q^{m-2}+\cdots+1}$
and $\operatorname{LT}(u_{m}d_{i,m})=(-1)^{\lfloor m/2\rfloor}\xi_{m-1}^{q^{m-1-i}+\cdots+1}\xi_{m}^{q^{m-1}+\cdots+q^{m-i}}$. Furthermore
$\nu(u_{m})=\nu(u_{m}d_{i,m})=1+q+\cdots+q^{m-1},$
$u_{m}=M(0,m)+\delta_{0,m}$ with $\delta_{0,m}\in R_{n-2}$, $u_{m}d_{i,m}=M(i,m)+\delta_{i,m}$ with $\delta_{i,m}\in R_{n-1}$ and, for $0\leq i\leq m$, $\nu(\delta_{i,m})>1+q+\cdots+q^{m-1}$.
(c) $\mathcal{H}$ is a homogeneous system of parameters.
(d) There exists $c_{n-1,m}\in R_{n-2}$ such that
$\xi_{n-1}=\left(\sum_{i=1}^{m}(-1)^{i+1}(\xi_{n-1-i}+c_{n-1-i,m})d_{i,m}\right)-c_{n-1,m}.$
(e) For $11\leq i\leq m-$ 1 there exist $\gamma_{k,m}^{(i)}\in R_{n-2}$ with $\nu(\gamma_{k,m}^{(i)})>q^{i}$ such that
$\xi_{n-1-i}^{q^{i}}=\left(\sum_{j=1}^{m}(-1)^{j+1}(\xi_{n-1-i-j}^{q^{i}}+\gamma_{n-1-i-j,m}^{(i)})d_{j,m}\right)-\gamma_{n-1-i,m}^{(i)}.$
@@
(f) $S_{m}^{G_{m}}$ is generated by $\mathcal{H}\cup\{\xi_{m},\ldots,\xi_{n-2}\}$. Furthermore, $S_{m}^{G_{m}}$ is the free $\mathbb{F}_{q}[\mathcal{H}]$-module with basis $\mathcal{B}$ and is a complete intersection with the relations given in (e).
@@
$\xi_{n-1-i}^{q^{i}}=\sum_{j=1}^{m}(-1)^{j+1}\xi_{n-1-i-j}^{q^{i}}d_{j,m}\quad\text{ for }i=1,2,\dots,m-1.$
@@
The invariants $\xi_{0},\xi_{1},\dots,\xi_{n-2}$ are algebraically independent and of degree less than $\deg(d_{i,m})$ for all $i$. Hence none of these $\xi_{i}$ is redundant. But Theorem 3.1 shows that at least $m$ other invariants are required to generate $S_{m}^{G_{m}}$. ∎

### v3

@@ 4. Formulating the Main Theorem
In the following, for $i<2m+1$, we identify $R_{i}$ and $\Phi_{i,m}(R_{i})$. Recall that $\psi:S_{m}\to S_{m}[t]$ is the algebra homomorphism determined by $\psi(a)=a^{q}-at^{q-1}$ for $a$ homogeneous of degree one. Note that $\psi$ followed by evaluating $t$ at $x_{m}$ gives $\psi_{[m,1]}:S_{m}\to S_{m}$. Using Corollary 3.6 and the relationship between $\psi$ and the complete Steenrod operator gives
(1) $\psi(\xi_{1})=\xi_{1}^{q}-\xi_{2}t^{q-1}+\xi_{1}t^{2(q-1)},$
(2) $\psi(\xi_{2})=\xi_{2}^{q}-\xi_{3}t^{q-1}-2\xi_{1}^{q}t^{q(q-1)}+\xi_{2}t^{(q+1)(q-1)}$
and, for $i\geq 3$,
(3) $\psi(\xi_{i})=\xi_{i}^{q}-\xi_{i+1}t^{q-1}-\xi_{i-1}^{q}t^{q^{i-1}(q-1)}+\xi_{i}t^{(q^{i-1}+1)(q-1)}\,.$
Therefore $\psi$ restricts to a map from $R_{n-1}$ to $R_{n}[t]$.
For $f\in R_{n}$, define ${R\mbox{-}\deg}(f)$ to be the degree of $f$ as a polynomial in the variables $\xi_{1},\ldots,\xi_{n}$ and define ${S\mbox{-}\deg}(f)$ to be the degree of $f$ as an element of $S_{m}$. For example, ${S\mbox{-}\deg}(\xi_{i})=q^{i-1}+1$ while ${R\mbox{-}\deg}(\xi_{i})=1$. We will use the weighted reverse lexicographic order on $R_{n}$ with $S\mbox{-}\deg$} as the weight. This means we first compare $S\mbox{-}\deg$} and if two monomials have the same $S\mbox{-}\deg$}, we use reverse lex to determine the order (see [18, Example 1.2.8]). In Magma [2], this order is called Graded Reverse Lexicographical (Weighted) and is implemented using the command grevlexw (see the online Magma handbook). Using this order $\xi_{i+1}>\xi_{i}$ and $\xi_{2}^{q+1}>\xi_{1}^{q}\xi_{3}>\xi_{2}^{q-1}\xi_{1}^{q+1}$ (all three monomials have $S\mbox{-}\deg$} equal to $1+2q+q^{2}$ and larger exponents on $\xi_{1}$ gives smaller monomials).
@@
Define $\nu:R_{n}\to\mathbb{Z}\cup\{\infty\}$ by $\nu(0)=\infty$ and, for $f\not=0$, $\nu(f)$ is the minimum value of $R\mbox{-}\deg$} on the non-zero terms of $f$. For example, $\nu(\xi_{2}^{q})=q$ and $\nu(\xi_{2}^{q}-\xi_{3}\xi_{2}^{(q-1)/2})=(q+1)/2$. Note that if $\nu(f)=\ell$, then $f\in(R_{n})_{+}^{\ell}\setminus(R_{n})_{+}^{\ell+1}$ where $(R_{n})_{+}$ denotes the ideal of $R_{n}$ generated by $\{\xi_{n},\xi_{n-1},\dots,\xi_{1}\}$. It is easy to verify that for $f,h\in R_{n}$, we have $\nu(fh)=\nu(f)\nu(h)$ and $\nu(f+h)\geq{\rm min}\{\nu(f),\nu(h)\}$. It follows from [30, Ch IV, §9, Theorem 14], that defining $\nu(f/h)=\nu(f)-\nu(h)$ extends $\nu$ to a discrete valuation on $\mathcal{F}(R_{n})$. Note that the associated Discrete Valuation Ring, $\{a\in\mathcal{F}(R_{n})\mid\nu(a)\geq 0\}$, properly contains $R_{n}$, i.e., $R_{n}$ is not the associated Discrete Valuation Ring.
@@
For $f\in R_{n}$, we have $\nu(\mathcal{P}^{i}(f))\geq\nu(f)$.
@@
Suppose $\beta$ is a monomial in $R_{n}$ with ${S\mbox{-}\deg}(\beta)={S\mbox{-}\deg}(\xi_{i}^{b}\xi_{i+1}^{a})$. If $\nu(\beta)>\nu(\xi_{i}^{b}\xi_{i+1}^{a})$ then $\operatorname{LT}(\beta)<\operatorname{LT}(\xi_{i}^{b}\xi_{i+1}^{a})$ using the weighted revlex order.
@@
It is sufficient to show that $\beta$ is in the ideal $\langle\xi_{1},\ldots,\xi_{i-1},\xi_{i}^{b+1}\rangle\subset R_{n}$. Suppose, by way of contradiction, that this is not the case. Since $\nu(\beta)>b+a$, $\beta$ has at least $a+b+1$ factors. Since $\beta$ is not in the ideal $\langle\xi_{1},\ldots,\xi_{i-1},\xi_{i}^{b+1}\rangle$, at least $a+1$ of these factors must be greater than $\xi_{i}$. The smallest weighted degree of such a monomial is
$b(q^{i-1}+1)+(a+1)(q^{i}+1)>b(q^{i-1}+1)+a(q^{i}+1)={S\mbox{-}\deg}(\xi_{i}^{b}\xi_{i+1}^{a}),$
@@
$M_{m}:=\begin{pmatrix}\xi_{n}&\xi_{n-1}&&\cdots&\xi_{m}\\ \xi_{n-1}^{q}&\xi_{n-2}^{q}&&\cdots&\xi_{m-1}^{q}\\ &&\vdots&&\\ \xi_{m+1}^{q^{m-1}}&\xi_{m}^{q^{m-1}}&&\cdots&\xi_{1}^{q^{m-1}}\end{pmatrix}$
and let $M(i,m)$ denote the minor formed by removing column $i+1$ from $M_{m}$. Note that $\nu(M(i,m))=1+q+\cdots+q^{m-1}$.
@@
Note that $\nu(m_{j,k})=q^{j-1}$. For $k>1$, it follows from Lemma 3.6 that $\mathcal{P}^{q^{n-k}}(m_{j,k})=m_{j,k-1}$, and $\nu(\mathcal{P}^{\ell}(m_{j,k}))>q^{j-1}$ unless $\ell=0$ or $\ell=q^{n-k}$. Therefore, since $e(i,m)=q^{n-2}+q^{n-3}+\cdots+q^{n-i-1}$, the Cartan identity yields
@@
Finally we are ready to formulate the main theorem. We will show that $S_{m}^{G_{m}}$ is generated by the 3m1=n+m13m-1=n+m-1 invariants
$\mathcal{S}_{m}:=\{\xi_{1},\ldots,\xi_{n-1},d_{1,m},\ldots,d_{m,m}\}.$
The proof is by induction on $m$ and relies on the following lemma.
Lemma 4.6.
Suppose $m>1$ and $S_{m-1}^{G_{m-1}}$ is generated by $\mathcal{S}_{m-1}$.
(a) $u_{m}\in R_{n-1}$, $\ker(\Phi_{n-1,m-1})=u_{m}R_{n-1}$, and $u_{m}d_{i,m}\in R_{n}$.
(b) $\mathcal{P}^{i}(u_{m})=0$ for $0<i<q^{m-1}$ and $\mathcal{P}^{i}(u_{m})/u_{m}\in R_{n-1}$ for $q^{m-1}\leq i<q^{n-2}$. For $i=\ell q^{n-2}+r$ with $0\leq\ell<q$ and $0\leq r<q^{n-2}$, $u_{m}^{\ell+1}\mathcal{P}^{i}(d_{j,m})\in R_{n}$ with $\xi_{n}$-degree at most $\ell+1$. If $\ell=q-1$ and $i\not\equiv_{(q)}0$ then $u_{m}^{q}\mathcal{P}^{i}(d_{j,m})$ has $\xi_{n}$-degree at most q1q-1.
(c) Interpreting $u_{m-1}$ and $u_{m-1}d_{i,m-1}$ as elements of $R_{n-2}\subset S_{m}^{G_{m}}$, we have
$u_{m}=(\xi_{n-1}+c_{n-1,m})u_{m-1}^{q}+\sum_{i=1}^{m-1}(-1)^{i}(\xi_{n-1-i}+c_{n-1-i,m})(u_{m-1}d_{i,m-1})^{q}$
with $c_{j,m}\in R_{j-1}$ and $u_{m}d_{i,m}-(u_{m-1}d_{i-1,m-1})^{q}\xi_{n}\in R_{n-1}$.
(d) Using the weighted revlex order on $R_{n}$,
$\operatorname{LT}(u_{m})=(-1)^{\lfloor m/2\rfloor}\xi_{m}^{q^{m-1}+q^{m-2}+\cdots+1}$
and $\operatorname{LT}(u_{m}d_{i,m})=(-1)^{\lfloor m/2\rfloor}\xi_{m}^{q^{m-1-i}+\cdots+1}\xi_{m+1}^{q^{m-1}+\cdots+q^{m-i}}$. Note that for $i=m$, $\operatorname{LT}(u_{m}d_{m,m})$ is a power of $\xi_{m+1}$.
Furthermore
$\nu(u_{m})=\nu(u_{m}d_{i,m})=1+q+\cdots+q^{m-1},$
$u_{m}=M(0,m)+\delta_{0,m}$ with $\delta_{0,m}\in R_{n-1}$, $u_{m}d_{i,m}=M(i,m)+\delta_{i,m}$ with $\delta_{i,m}\in R_{n}$ and, for $0\leq i\leq m$, $\nu(\delta_{i,m})>1+q+\cdots+q^{m-1}$.
(e) There exists $c_{n,m}\in R_{n-1}$ such that
$\xi_{n}=\left(\sum_{i=1}^{m}(-1)^{i+1}(\xi_{n-i}+c_{n-i,m})d_{i,m}\right)-c_{n,m}.$
(f) For $1\leq i\leq m-1$ there exist $\gamma_{k,m}^{(i)}\in R_{n-1}$ with $\nu(\gamma_{k,m}^{(i)})>q^{i}$ such that
$\xi_{n-i}^{q^{i}}=\left(\sum_{j=1}^{m}(-1)^{j+1}(\xi_{n-i-j}^{q^{i}}+\gamma_{n-i-j,m}^{(i)})d_{j,m}\right)-\gamma_{n-i,m}^{(i)}.$
We postpone the proof of Lemma 4.6 to Section 7.
Remark 4.7.
It follows from Lemma 4.2(b) and Lemma 4.6(a) that if $u_{m}$ divides $f$ in $S_{m}$ and $f\in R_{n-1}$ then $f/u_{m}\in R_{n-1}$. To see this, suppose $u_{m}$ divides $f$ in $S_{m}$ and $f\in R_{n-1}$. Then $x_{m}$ divides $f$ in $S_{m}$, which means that $f\in\ker(\psi_{[m,1]})$. Since $f\in R_{n-1}$, we have $f\in\ker(\psi_{[m,1]}\circ\Phi_{n-1,m})$. Using Lemma 4.2(b) gives $\ker(\Phi_{n-1,m-1})=\ker(\psi_{[m,1]}\circ\Phi_{n-1,m})$. From Lemma 4.6(a), we have $\ker(\Phi_{n-1,m-1})=u_{m}R_{n-1}$. Therefore $f\in u_{m}R_{n-1}$, as required.
Recall from Theorem 3.12 the homogeneous system of parameters
$\mathcal{H}=\{\xi_{1},\ldots,\xi_{m},d_{1,m},\ldots,d_{m,m}\}$
where $\deg(\xi_{i})=q^{i-1}+1$ and $\deg(d_{i,m})=q^{n-1}-q^{n-1-i}$. Let $\mathcal{B}$ denote the set of monomial factors of
$\Gamma=\prod_{i=m+1}^{n-1}\xi_{i}^{q^{n-i}-1}=\prod_{j=1}^{m-1}\xi_{n-j}^{q^{j}-1}\ .$
We now state our main theorem.
@@
Suppose $m>1$. $S_{m}^{G_{m}}$ is generated by
$\mathcal{S}_{m}=\mathcal{H}\cup\{\xi_{m+1},\ldots,\xi_{n-1}\}=\{\xi_{1},\ldots,\xi_{n-1},d_{1,m},\ldots,d_{m,m}\}.$
@@
Furthermore, $S_{m}^{G_{m}}$ is the free $\mathbb{F}_{q}[\mathcal{H}]$-module with basis $\mathcal{B}$ and is the complete intersection with the relations given in part (f) of Lemma 4.6.
Proof.
We assume the conclusions of Lemma 4.6 hold for $S_{m}^{G_{m}}$. Note that $S_{1}^{G_{1}}$ is the polynomial algebra generated by $\xi_{1}=x_{1}y_{1}$ and $\xi_{2}/\xi_{1}=y_{1}^{q-1}+x_{1}^{q-1}$.
Let $A$ denote the subalgebra of $S_{m}^{G_{m}}$ generated by $\mathcal{S}_{m}$. Using Theorem 3.12, $A\subseteq S_{m}^{G_{m}}$ is an integral extension. Using part (e) of Lemma 4.6, $R_{n}\subset A$. Therefore $\mathcal{F}(A)=\mathcal{F}(R_{n})=\mathcal{F}(S_{m}^{G_{m}})$. Thus to show $A=S_{m}^{G_{m}}$, it is sufficient to show that $A$ is integrally closed in its field of fractions. We will use [19, Proposition 1.1] to prove that $A$ is a UFD. Recall that a UFD is integrally closed in its field of fractions (see, for example, [13, Proposition 4.10]).
Using the definition and part (f) of Lemma 4.6, $A$ is the $\mathbb{F}_{q}[\mathcal{H}]$-module generated by $\mathcal{B}$. Note that the number of elements of the set $\mathcal{B}$ is $\prod_{i=1}^{m-1}q^{i}=q^{m(m-1)/2}$. From Theorem 5.11, we know that $S_{m}^{P_{m}}$ is Cohen-Macaulay. Therefore $S_{m}^{G_{m}}$ is Cohen-Macaulay and $S_{m}^{G_{m}}$ is a free $\mathbb{F}_{q}[\mathcal{H}]$-module of rank
which simplifies to $q^{m(m-1)/2}$. Therefore the field extension $\mathcal{F}(\mathbb{F}_{q}[\mathcal{H}])\subset\mathcal{F}(S_{m}^{G_{m}})$ is of degree $q^{m(m-1)/2}$. Since $\mathcal{F}(A)=\mathcal{F}(S_{m}^{G_{m}})$, the degree of the field extension $\mathcal{F}(\mathbb{F}_{q}[\mathcal{H}])\subset\mathcal{F}(A)$ is $q^{m(m-1)/2}$. We know that $\mathcal{B}$ is a spanning set for $\mathcal{F}(A)$ over $\mathcal{F}(\mathbb{F}_{q}[\mathcal{H}])$. By comparing the order of $\mathcal{B}$ with the degree of the extension, we see that $\mathcal{B}$ is linearly independent over $\mathcal{F}(\mathbb{F}_{q}[\mathcal{H}])$. Therefore $A$ is a free $\mathbb{F}_{q}[\mathcal{H}]$-module with basis $\mathcal{B}$ and $A$ is Cohen-Macaulay. Since $A$ is a free $\mathbb{F}_{q}[\mathcal{H}]$-module with basis $\mathcal{B}$, the Hilbert series for $A$ is
$HS(A,t)=\left(\prod_{i=1}^{m}\frac{1}{1-t^{q^{n-1}-q^{n-1-i}}}\right)\left(\prod_{i=1}^{m}\frac{1}{1-t^{q^{i-1}+1}}\right)\left(\prod_{i=m+1}^{n-1}\frac{1-t^{q^{n-1}+q^{n-i}}}{1-t^{q^{i-1}+1}}\right).$
The complete intersection with relations given by Lemma 4.6(f) surjects onto $A$. Since the the Hilbert series of the complete intersection coincides with the Hilbert series for $A$, the surjection is a bijection. Hence $A$ is the complete intersection with relations given in Lemma 4.6(f). Therefore, to complete the proof of the theorem, we need only show that $A=S_{m}^{G_{m}}$.
From Lemma 4.6(d), for $m$ and m1m-1, we have
$\operatorname{LT}(u_{m})=(-1)^{\lfloor m/2\rfloor}\xi_{m}^{q^{m-1}+q^{m-2}+\cdots+1}$
and $\operatorname{LT}(u_{m-1})=(-1)^{\lfloor(m-1)/2\rfloor}\xi_{m-1}^{q^{m-2}+q^{m-3}+\cdots+1}$ (for $m=2$, $u_{m-1}=u_{1}=x_{1}y_{1}=\xi_{1}$). Therefore, since $\mathcal{H}$ is an homogeneous system of parameters, $\{u_{m-1},u_{m}\}$ is a partial homogeneous system of parameters. Since $A$ is Cohen-Macaulay, this means that $u_{m-1},u_{m}]$ is a regular sequence in $A$. From Lemma 4.6(a) for m1m-1, when $m>2$, we have $\ker(\Phi_{n-3,m-2})=u_{m-1}R_{n-3}$. Therefore $u_{m-1}$ is prime in $R_{n-3}$. If $m=2$, $u_{1}=\xi_{1}$ is prime in $R_{1}=\mathbb{F}_{q}[\xi_{1}]$. Since $R_{n}$ is a polynomial extension of $R_{n-3}$, $u_{m-1}$ is prime in $R_{n}$ and in its localisation $R_{n}[u_{m}^{-1}]$. Using part (e) of Lemma 4.6 , $R_{n}\subset A$ and by definition $u_{m}d_{i,m}\in R_{n}$. Therefore $A[u_{m}^{-1}]=R_{n}[u_{m}^{-1}]$ and $u_{m-1}$ is prime in $A[u_{m}^{-1}]$. To satisfy the hypotheses of [19, Proposition 1.1] and complete the proof, we need to show that $A[u_{m-1}^{-1}]$ is a UFD.
From Lemma 4.6(c), we have $u_{m}d_{1,m}-u_{m-1}^{q}\xi_{n}\in R_{n-1}$. Therefore $R_{n}\subset R_{n-1}[d_{1,m},u_{m-1}^{-1}]$ and $\mathcal{F}(R_{n})=\mathcal{F}(R_{n-1}[d_{1,m}])$. Hence $R_{n-1}[d_{1,m}]$ is a polynomial algebra and $R_{n-1}[d_{1,m},u_{m-1}^{-1}]$ is a UFD. We prove $A[u_{m-1}^{-1}]$ is a UFD by showing $A[u_{m-1}^{-1}]=R_{n-1}[d_{1,m},u_{m-1}^{-1}]$. By definition, $R_{n-1}[d_{1,m}]\subset A$. From Lemma 4.6(c), we have $u_{m}d_{i,m}-(u_{m-1}d_{i-1,m-1})^{q}\xi_{n}\in R_{n-1}$ for $i>1$. Cross-multiplying to eliminate $\xi_{n}$ gives
$u_{m-1}^{q}u_{m}d_{i,m}-(u_{m-1}d_{i-1,m-1})^{q}u_{m}d_{1,m}\in R_{n-1}.$
Using $\ker(\Phi_{n-1,m-1})=u_{m}R_{n-1}$ (Lemma 4.6(a)) and Remark 4.7, we see that
$u_{m-1}^{q}d_{i,m}-(u_{m-1}d_{i-1,m-1})^{q}d_{1,m}\in R_{n-1}$
(with $u_{m-1}d_{i-1,m-1}\in R_{n-2}$). Thus $d_{i,m}\in R_{n-1}[d_{1,m},u_{m-1}^{-1}]$, as required. ∎
@@
$\xi_{n-i}^{q^{i}}=\sum_{j=1}^{m}(-1)^{j+1}\xi_{n-i-j}^{q^{i}}d_{j,m}\quad\text{ for }i=1,2,\dots,m-1.$
@@
The invariants $\xi_{1},\xi_{2},\dots,\xi_{n-1}$ are algebraically independent and of degree less than $\deg(d_{i,m})$ for all $i$. Hence none of these $\xi_{i}$ is redundant. But Theorem 3.1 shows that at least $m$ other invariants are required to generate $S_{m}^{G_{m}}$. ∎